Two coils connected in series have resistance of 600ohm and 300ohm at 20°c and temperature co-efficient of 0.001and 0.004(°C)^-1 respectively. Find the resistance of the combination at a temperature of 50°C. What is the effective temperature co-efficient of combination? Ans-954ohm ,0.002 /°c how??????explain


Grade:10

1 Answers

Sanjhi Priya
4 years ago
R1' = R1'(1+a1*∆T)  = 600(1+0.001*(50-20))= 618 ohm
Similarly R2 = 300(1+0.004*(50-20)) = 336 ohm
Hence, Req. = 618+336= 954 ohm
Req' = Req (1+aeq.*∆T) which gives aeq. =
(954/900)-1/30 = 1/500